A bar magnet has a length of 8 cm. The magnetic field at a point at a distance 3 cm from the centre in the broadside-on position is found to be 4 × 10^-6 T. Find the pole strength of the magnet.

Grade:11

1 Answers

Kevin Nash
9 years ago
Sol. Magnetic field at the broad side on position : B = μ base 0/4π M/(d^2 + ℓ^2)^3/2 2ℓ = 8 cm d = 3 cm ⇒ 4 * 10^-6 = 10^-7 *m * 8 * 10^-2/(9 * 10^-4 + 16 *10^-4)^3/2 ⇒ 4 * 10^-6 = 10^-9 *m * 8/(10^-4)^3/2 +(25)3/2 ⇒ m = 4 * 10^-6 * 125 * 10^-8/8 * 10^-9 = 62.5 * 10^-5 A-m